A ramp 17.5 feet in the length rises to a loading platform that is 3 1/3 feet off the ground. Find the angle that the ramp makes with the ground?
Answer by scott8148(5880)

the sine of the angle is the height of the platform divided by the length of the ramp

let x=angle __ sin(x)=(10/3)/(35/2) __ sin(x)=4/21

use a calculator (or a table) to find x